There are 400kf/min of water being handled by a pump. The lift us firm 20-m deep well and the delivery velocity ks 15m/s. Find (a) the change in potential energy, (b) kinetic energy, (c) the required power of the pumping unit; g= 9.81m/s²

1
Expert's answer
2022-04-24T17:18:29-0400

a)

"S=\frac Qv=0 .44\cdot 10^{-3}~m^2,"

"V=Sh=8.9\cdot 10^{-3}~m^3,"

"m=\rho V=8.9~kg,"

"U=mgh=1780~J,"

b)

"E=\frac{mv^2}2=1000~J,"

c)

"P=mgv=133~W."